# Characteristic Learning for Provable One Step Generation

### v3

@@ C.1. Auxiliary Properties
=$\stackrel{{\scriptstyle\,\mathrm{d}}}{{=}}\frac{\alpha_{t}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}(U-{\mathbb{E}}[U])+\sqrt{\frac{\sigma^{2}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}}\epsilon.$
Since that the random variable $U\in[0,1]^{d}$, using Hoeffding’s lemma [Mohri et al., 2018, Lemma D.1] implies that $U$ is $1$-sub-Gaussian. Further, applying Lemma A.4 deduces that each element of $(\varepsilon_{t}|X_{t}=x)$ is sub-Gaussian for each $t\in[0,T]$ and $x\in{\mathbb{R}}^{d}$ with variance proxy
(D.3) $\sigma_{T}^{2}=\sup_{t\in[0,T]}\Big{\{}\frac{\alpha_{t}^{2}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})^{2}}{(\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2})^{2}}+\frac{\sigma^{2}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\Big{\}}\leq C\lambda(T),$
@@
^$\widehat{{\mathcal{E}}}_{T,R,n}(b)$ $=\frac{1}{n}\sum_{i=1}^{n}\|b^{*}(t^{(i)},X_{t}^{(i)})-b(t^{(i)},X_{t}^{(i)})\|_{2}^{2}\mathbbm{1}\{\|X_{t}^{(i)}\|_{\infty}\leq R\}.$
@@
(D.4) ${\mathbb{E}}_{{\mathcal{S}}}[{\mathcal{E}}_{T}(\widehat{b})]\leq{\mathbb{E}}_{{\mathcal{S}}}\Big{[}{\mathcal{E}}_{T}(\widehat{b})-{\mathcal{E}}_{T,R}(\widehat{b})\Big{]}+{\mathbb{E}}_{{\mathcal{S}}}\Big{[}\sup_{b\in{\mathscr{B}}}{\mathcal{E}}_{T,R}(b)-2\widehat{{\mathcal{E}}}_{T,R,n}(b)\Big{]}+2{\mathbb{E}}_{{\mathcal{S}}}\Big{[}\widehat{{\mathcal{E}}}_{T,R,n}(\widehat{b})\Big{]}.$
The first term in the right-hand side of (D.4) corresponds to the truncation error, which is estimated in the rest of this step. The second term of (D.4) is studied in Step 3. Finally, we bound the last term of (D.4) in Step 4.
@@
(D.5) $\leq 8\Big{(}{\mathbb{E}}_{X_{t}\sim\mu_{t}}^{1/2}\Big{[}\|b(t,X_{t})\|_{2}^{4}\Big{]}+{\mathbb{E}}_{X_{t}\sim\mu_{t}}^{1/2}\Big{[}\|b^{*}(t,X_{t})\|_{2}^{4}\Big{]}\Big{)}\operatorname{Pr}^{1/2}\{\|X_{t}\|_{\infty}>R\},$
where the first inequality follows from Cauchy-Schwarz inequality, and the second inequality is due to the triangular inequality. The boundedness of the hypothesis (D.1) deduces
(D.6) ${\mathbb{E}}_{X_{t}\sim\mu_{t}}^{1/2}\Big{[}\|b(t,X_{t})\|_{2}^{4}\Big{]}\leq dB_{\operatorname{vel}}^{2}R^{2}.$
Then we consider the fourth moment of $b^{*}(t,X_{t})$ in (D.5). By using Assumption 2, we have
$Y_{t}=\dot{\alpha}_{t}X_{0}+\dot{\beta}_{t}X_{1}\stackrel{{\scriptstyle\,\mathrm{d}}}{{=}}\dot{\alpha}_{t}X_{0}+\dot{\beta}_{t}U+\sigma\dot{\beta}_{t}\epsilon,\quad U\sim\nu,\leavevmode\nobreak\ \epsilon\sim N(0,I_{d}),$
@@
where the second inequality holds from the triangular inequality, the last inequality follows from Lemma A.7, and the constant $C$ only depends in $d$ and $\sigma$. Consequently,
(D.7) ${\mathbb{E}}_{X_{t}}^{1/2}\Big{[}\|b^{*}(t,X_{t})\|_{2}^{4}\Big{]}={\mathbb{E}}_{X_{t}}^{1/2}\Big{[}\|{\mathbb{E}}[Y_{t}|X_{t}]\|_{2}^{4}\Big{]}\leq{\mathbb{E}}^{1/2}\Big{[}\|Y_{t}\|_{2}^{4}\Big{]}\leq C\lambda(T),$
where the first inequality is due to the definition of velocity, and the second inequality follows from Jensen’s inequality. We next consider the tail probability of $X_{t}$ in (D.5). According to Assumption 2, we find
$X_{t}\stackrel{{\scriptstyle\,\mathrm{d}}}{{=}}\alpha_{t}X_{0}+\beta_{t}U+\sigma\beta_{t}\epsilon,\quad U\sim\nu,\leavevmode\nobreak\ \epsilon\sim N(0,I_{d}),$
which implies from Hoeffding’s lemma [Mohri et al., 2018, Lemma D.1] and Lemma A.4 that $X_{t}$ is a $(\alpha_{t}^{2}+\beta_{t}^{2}+\sigma^{2}\beta_{t}^{2})$-sub-Gaussian random variable. Then it follows from Lemma A.5 that
(D.8) $\sup_{t\in(0,1)}\operatorname{Pr}\big{\{}\|X_{t}\|_{\infty}>R\big{\}}\leq 2d\sup_{t\in(0,1)}\exp\Big{(}-\frac{R^{2}}{2(\alpha_{t}^{2}+\beta_{t}^{2}+\sigma^{2}\beta_{t}^{2})}\Big{)}\leq\frac{2d}{\exp(\theta R^{2})},$
where $\theta$ is a constant only depending on $\sigma$. Substituting (D.6), (D.7), and (D.8) into (D.5) yields
@@
(D.9) ${\mathcal{E}}_{T}(b)-{\mathcal{E}}_{T,R}(b)\leq\frac{C\lambda(T)R^{2}}{\exp(\theta R^{2})}.$
Hence the first term in the right-hand side of (D.4) can be bounded by
(D.10) ${\mathbb{E}}_{{\mathcal{S}}}\Big{[}{\mathcal{E}}_{T}(\widehat{b})-{\mathcal{E}}_{T,R}(\widehat{b})\Big{]}\leq\frac{C\lambda(T)R^{2}}{\exp(\theta R^{2})}.$
We next bound another truncation term by a similar argument, which will be used in Step 4. It follows from Cauchy-Schwarz inequality and (D.8) that
(D.11) ${\mathbb{E}}_{{\mathcal{S}}}\Big{[}(\varepsilon_{t,k}^{(i)})^{2}\mathbbm{1}\{\|X_{t}^{(i)}\|_{\infty}>R\}\Big{]}\leq{\mathbb{E}}_{{\mathcal{S}}}^{1/2}\Big{[}(\varepsilon_{t,k}^{(i)})^{4}\Big{]}\operatorname{Pr}^{1/2}\{\|X_{t}^{(i)}\|_{\infty}>R\}\leq\frac{C\lambda(T)R^{2}}{\exp(\theta R^{2})},$
where $C$ is a constant only depending on $d$ and $\sigma$, and the last inequality follows from the fact that $\varepsilon_{t,k}^{(i)}$ is sub-Gaussian with variance proxy in (D.3).
@@
(D.12) ${\mathbb{E}}_{{\mathcal{S}}}\Big{[}\sup_{b\in{\mathscr{B}}}{\mathcal{E}}_{T,R}(b)-2\widehat{{\mathcal{E}}}_{T,R,n}(b)\Big{]}\leq CR^{2}\sum_{k=1}^{d}\frac{\operatorname{VCdim}(\Pi_{k}{\mathscr{B}})}{n\log^{-1}(n)},$
@@
^$\widehat{{\mathcal{E}}}_{T,R,n}^{k}(b)$ $=\frac{1}{n}\sum_{i=1}^{n}(b_{k}^{*}(t^{(i)},X_{t}^{(i)})-b_{k}(t^{(i)},X_{t}^{(i)}))^{2}\mathbbm{1}\{\|X_{t}^{(i)}\|_{\infty}\leq R\}.$
Applying Proposition 3.2, (D.1), and Lemma G.1 yields the following inequality
@@
which implies (D.12) by summing with respect to $1\leq k\leq d$.
@@
(D.13) ${\mathbb{E}}_{{\mathcal{S}}}\Big{[}\widehat{{\mathcal{E}}}_{T,R,n}(\widehat{b})\Big{]}\leq 2\inf_{b\in{\mathscr{B}}}{\mathcal{E}}_{T,R}(b)+C\lambda(T)R^{2}\Big{(}\sum_{k=1}^{d}\frac{\operatorname{VCdim}(\Pi_{k}{\mathscr{B}})}{n\log^{-1}n}+\frac{1}{\exp(\theta R^{2})}\Big{)},$
@@
^$\widehat{{\mathcal{L}}}_{T,R,n}(\widehat{b})$ $=\frac{1}{n}\sum_{i=1}^{n}\sum_{k=1}^{d}(\varepsilon_{t,k}^{(i)}+b_{k}^{*}(t^{(i)},X_{t}^{(i)})-\widehat{b}_{k}(t^{(i)},X_{t}^{(i)}))^{2}\mathbbm{1}\{\|X_{t}^{(i)}\|_{\infty}\leq R\}$
=^$=\widehat{{\mathcal{E}}}_{T,R,n}(\widehat{b})+\frac{1}{n}\sum_{i=1}^{n}\sum_{k=1}^{d}(\varepsilon_{t,k}^{(i)})^{2}\mathbbm{1}\{\|X_{t}^{(i)}\|_{\infty}\leq R\}$
@@
Since $\widehat{b}$ is a minimizer of ^$\widehat{{\mathcal{L}}}_{T,n}(\cdot)$ over the hypothesis class ${\mathscr{B}}$, it holds that ^$\widehat{{\mathcal{L}}}_{T,R,n}(\widehat{b})\leq\widehat{{\mathcal{L}}}_{T,n}(\widehat{b})\leq\widehat{{\mathcal{L}}}_{T,n}(b)$ for each $b\in{\mathscr{B}}$. Consequently,
(D.14) ${\mathbb{E}}_{{\mathcal{S}}}\Big{[}\widehat{{\mathcal{E}}}_{T,R,n}(\widehat{b})\Big{]}$ $\leq{\mathcal{L}}_{T}(b)-{\mathbb{E}}_{{\mathcal{S}}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\sum_{k=1}^{d}(\varepsilon_{t,k}^{(i)})^{2}\mathbbm{1}\{\|X_{t}^{(i)}\|_{\infty}\leq R\}\Big{]}$
@@
where we used the fact that ${\mathbb{E}}[\widehat{{\mathcal{L}}}_{T,n}(b)]={\mathcal{L}}_{T}(b)$ and ${\mathbb{E}}[\varepsilon_{t,k}^{(i)}b^{*}(t^{(i)},X_{t}^{(i)})]=0$. For the first term in the right-hand side of (D.14), we have
@@
(D.15) $={\mathcal{E}}_{T}(b)+{\mathbb{E}}_{{\mathcal{S}}}\Big{[}\frac{1}{n}\sum_{i=1}^{n}\sum_{k=1}^{d}(\varepsilon_{t,k}^{(i)})^{2}\mathbbm{1}\{\|X_{t}^{(i)}\|_{\infty}>R\}\Big{]}\leq{\mathcal{E}}_{T,R}(b)+2\frac{C\lambda(T)R^{2}}{\exp(\theta R^{2})},$
where the inequality follows from (D.9) and (D.11). It remains to bound the second term in the right-hand side of (D.14). Plugging Proposition 3.2 and (D.3) into Lemma G.2, we have
@@
(D.16) $\leq\frac{1}{4}{\mathbb{E}}_{{\mathcal{S}}}\Big{[}\widehat{{\mathcal{E}}}_{T,R,n}(\widehat{b})\Big{]}+C\lambda(T)R^{2}\sum_{k=1}^{d}\frac{\operatorname{VCdim}(\Pi_{k}{\mathscr{B}})}{n\log^{-1}n},$
where $C$ is a constant only depending $d$ and $\sigma$. Substituting (D.15) and (D.16) into (D.14) yields (D.13).
Finally, plugging (D.10), (D.12), and (D.13) into (D.4) completes the proof. ∎
Lemma D.2 (Approximation error).
@@
=${\mathscr{B}}=\left\{b\in N(L,S):\begin{aligned} &\|b(t,x)\|_{\infty}\leq B_{\operatorname{vel}}R,\leavevmode\nobreak\ \|\partial_{t}b(t,x)\|_{2}\leq 3D\kappa(T)R,\\ &\|\nabla b(t,x)\|_{\operatorname{op}}\leq 3G,\leavevmode\nobreak\ (t,x)\in[0,T]\times{\mathbb{R}}^{d}\end{aligned}\right\},$
@@
Proof of Lemma D.2.
Denote by ${\mathbb{B}}_{R,T}^{\infty}=[0,T]\times{\mathbb{B}}_{R}^{\infty}$. According to Corollary H.6, for each element $1\leq k\leq d$, there exists a real-valued deep neural network $b_{k}$ with depth $L_{k}=\lceil\log_{2}(d+1)\rceil+3$ and number of parameters $S_{k}=(22(d+1)+6)(N+1)^{d+1}$, such that
@@
Set the hypothesis class as that in Lemma D.2. According to Lemma D.2, there exists a vector-valued deep neural network $b\in{\mathscr{B}}$ such that
(D.17) $\inf_{b\in{\mathscr{B}}}{\mathcal{E}}_{T,R}(b)\leq C\kappa^{2}(T)R^{2}N^{-2},$
where $C$ is a constant only depending on $d$ and $\sigma$. On the other hand, by applying Lemma A.12, the VC-dimension of this deep neural network class ${\mathscr{B}}$ is given as
(D.18) $\operatorname{VCdim}(\Pi_{k}{\mathscr{B}})\leq CN^{d+1}\log N,$
where $C$ is an absolute constant. Plugging (D.17) and (D.18) into Lemma D.1 yields
${\mathbb{E}}_{{\mathcal{S}}}\big{[}{\mathcal{E}}_{T}(\widehat{b})\big{]}$ $\leq CR^{2}\Big{(}\frac{\kappa^{2}(T)}{N^{2}}+\frac{\lambda(T)N^{d+1}\log N}{n\log^{-1}n}+\frac{\lambda(T)}{\exp(\theta R^{2})}\Big{)}$
@@
(D.19) ${\mathbb{E}}_{{\mathcal{S}}}\big{[}{\mathcal{E}}_{T}(\widehat{b})\big{]}\leq C\kappa^{2}(T)R^{2}\Big{(}n^{-\frac{2}{d+3}}\log(n)+\exp(-\theta R^{2})\Big{)}$
@@
(D.20) $\,\mathrm{d}Z(t)$ $=b^{*}(t,Z(t))\,\mathrm{d}t,\quad t\in(0,T),$
@@
(D.21) $\,\mathrm{d}\widehat{Z}(t)$ $=\widehat{b}(t,\widehat{Z}(t))\,\mathrm{d}t,\quad t\in(0,T),$
@@
(D.22) $\leq 2\|Z(t)-\widehat{Z}(t)\|_{2}\|b^{*}(t,Z(t))-\widehat{b}(t,\widehat{Z}(t))\|_{2},$
where the second inequality follows from (D.20) and (D.21), and the inequality holds from Cauchy-Schwarz inequality. On the other hand, we find
(D.23) $\frac{\,\mathrm{d}}{\,\mathrm{d}t}\|Z(t)-\widehat{Z}(t)\|_{2}^{2}=2\|Z(t)-\widehat{Z}(t)\|_{2}\frac{\,\mathrm{d}}{\,\mathrm{d}t}\|Z(t)-\widehat{Z}(t)\|_{2}.$
Combining (D.22) and (D.23) implies
@@
Then we turn to estimate the distribution error using the particle error bound derived in Lemma D.3.
@@
According to Lemma D.3, we have
@@
where the first inequality follows from the definition of $2$-Wasserstein distance and Jensen’s inequality. Substituting Theorem 3.9 into the above inequality completes the proof. ∎
Appendix E Proofs of Results in Section 3.4
@@
(E.1) $\widehat{Z}_{k}$ $=\widehat{Z}_{k-1}+\widehat{b}(t_{k-1},\widehat{Z}_{k-1})\tau,\quad 1\leq k\leq K,$
@@
Lemma E.1 (Discretization error of Euler method).
@@
(E.2) $=\widehat{Z}(t_{k})+\widehat{b}(t_{k},\widehat{Z}(t_{k}))\tau+\partial_{t}\widehat{b}(\theta,\widehat{Z}(\theta))\tau^{2},$
@@
(E.3) $\widehat{Z}_{k+1}=\widehat{Z}_{k}+\widehat{b}(t_{k},\widehat{Z}_{k})\tau.$
Subtracting (E.3) from (E.2) yields
@@
(E.4) $\leq(1+\|{\mathscr{B}}\|_{\operatorname{Lip}}\tau)\|\widehat{Z}(t_{k})-\widehat{Z}_{k}\|_{2}+D\kappa(T)R\tau^{2},$

### v5

@@ D.1. Auxiliary Properties
$\stackrel{{\scriptstyle\,\mathrm{d}}}{{=}}\frac{\alpha_{t}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}(U_{t,x}-{\mathbb{E}}[U_{t,x}])+\sqrt{\frac{\sigma^{2}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}}\epsilon.$
Since that the random variable $U_{t,x}\in[0,1]^{d}$, using Hoeffding’s lemma [Mohri et al., 2018, Lemma D.1] implies that $U_{t,x}$ is $1$-sub-Gaussian. Further, applying Lemma B.4 deduces that each element of $(\varepsilon_{t}|X_{t}=x)$ is sub-Gaussian for each $t\in[0,T]$ and $x\in{\mathbb{R}}^{d}$ with variance proxy
(E.3) $\sigma_{T}^{2}=\sup_{t\in[0,T]}\Big\{\frac{\alpha_{t}^{2}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})^{2}}{(\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2})^{2}}+\frac{\sigma^{2}(\alpha_{t}\dot{\beta}_{t}-\dot{\alpha}_{t}\beta_{t})^{2}}{\alpha_{t}^{2}+\sigma^{2}\beta_{t}^{2}}\Big\}\leq C\lambda(T),$
@@
$\widehat{{\mathcal{E}}}_{T,R,n}(b)$ $=\frac{1}{n}\sum_{i=1}^{n}\|b^{*}(t^{(i)},X_{t}^{(i)})-b(t^{(i)},X_{t}^{(i)})\|_{2}^{2}\mathbbm{1}\{\|X_{t}^{(i)}\|_{\infty}\leq R\}.$
@@
[…]
The first term in the right-hand side of (E.4) corresponds to the truncation error, which is estimated in the rest of this step. The second term of (E.4) is studied in Step 3. Finally, we bound the last term of (E.4) in Step 4.
@@
(E.5) $\leq 8\Big({\mathbb{E}}_{X_{t}\sim\mu_{t}}^{1/2}\Big[\|b(t,X_{t})\|_{2}^{4}\Big]+{\mathbb{E}}_{X_{t}\sim\mu_{t}}^{1/2}\Big[\|b^{*}(t,X_{t})\|_{2}^{4}\Big]\Big)\operatorname{Pr}^{1/2}\{\|X_{t}\|_{\infty}>R\},$
where the first inequality follows from Cauchy-Schwarz inequality, and the second inequality is due to the triangular inequality. The boundedness of the hypothesis (E.1) deduces
(E.6) ${\mathbb{E}}_{X_{t}\sim\mu_{t}}^{1/2}\Big[\|b(t,X_{t})\|_{2}^{4}\Big]\leq dB_{\operatorname{vel}}^{2}R^{2}.$
Then we consider the fourth moment of $b^{*}(t,X_{t})$ in (E.5). By using Assumption 2, we have
$Y_{t}=\dot{\alpha}_{t}X_{0}+\dot{\beta}_{t}X_{1}\stackrel{{\scriptstyle\,\mathrm{d}}}{{=}}\dot{\alpha}_{t}X_{0}+\dot{\beta}_{t}U+\sigma\dot{\beta}_{t}\epsilon,\quad U\sim\nu,\penalty 10000\ \epsilon\sim N(0,I_{d}),$
@@
where the second inequality holds from the triangular inequality, the last inequality follows from Lemma B.7, and the constant $C$ only depends in $d$ and $\sigma$. Consequently,
(E.7) ${\mathbb{E}}_{X_{t}}^{1/2}\Big[\|b^{*}(t,X_{t})\|_{2}^{4}\Big]={\mathbb{E}}_{X_{t}}^{1/2}\Big[\|{\mathbb{E}}[Y_{t}|X_{t}]\|_{2}^{4}\Big]\leq{\mathbb{E}}^{1/2}\Big[\|Y_{t}\|_{2}^{4}\Big]\leq C\lambda(T),$
where the first inequality is due to the definition of velocity, and the second inequality follows from Jensen’s inequality. We next consider the tail probability of $X_{t}$ in (E.5). According to Assumption 2, we find
$X_{t}\stackrel{{\scriptstyle\,\mathrm{d}}}{{=}}\alpha_{t}X_{0}+\beta_{t}U+\sigma\beta_{t}\epsilon,\quad U\sim\nu_{t,x},\penalty 10000\ \epsilon\sim N(0,I_{d}),$
which implies from Hoeffding’s lemma [Mohri et al., 2018, Lemma D.1] and Lemma B.4 that $X_{t}$ is a $(\alpha_{t}^{2}+\beta_{t}^{2}+\sigma^{2}\beta_{t}^{2})$-sub-Gaussian random variable. Then it follows from Lemma B.5 that
(E.8) $\sup_{t\in(0,1)}\operatorname{Pr}\big\{\|X_{t}\|_{\infty}>R\big\}\leq 2d\sup_{t\in(0,1)}\exp\Big(-\frac{R^{2}}{2(\alpha_{t}^{2}+\beta_{t}^{2}+\sigma^{2}\beta_{t}^{2})}\Big)\leq\frac{2d}{\exp(\theta R^{2})},$
where $\theta$ is a constant only depending on $\sigma$. Substituting (E.6), (E.7), and (E.8) into (E.5) yields
@@
(E.9) ${\mathcal{E}}_{T}(b)-{\mathcal{E}}_{T,R}(b)\leq\frac{C\lambda(T)R^{2}}{\exp(\theta R^{2})}.$
Hence the first term in the right-hand side of (E.4) can be bounded by
(E.10) ${\mathbb{E}}_{{\mathcal{S}}}\Big[{\mathcal{E}}_{T}(\widehat{b})-{\mathcal{E}}_{T,R}(\widehat{b})\Big]\leq\frac{C\lambda(T)R^{2}}{\exp(\theta R^{2})}.$
We next bound another truncation term by a similar argument, which will be used in Step 4. It follows from Cauchy-Schwarz inequality and (E.8) that
(E.11) ${\mathbb{E}}_{{\mathcal{S}}}\Big[(\varepsilon_{t,k}^{(i)})^{2}\mathbbm{1}\{\|X_{t}^{(i)}\|_{\infty}>R\}\Big]\leq{\mathbb{E}}_{{\mathcal{S}}}^{1/2}\Big[(\varepsilon_{t,k}^{(i)})^{4}\Big]\operatorname{Pr}^{1/2}\{\|X_{t}^{(i)}\|_{\infty}>R\}\leq\frac{C\lambda(T)R^{2}}{\exp(\theta R^{2})},$
where $C$ is a constant only depending on $d$ and $\sigma$, and the last inequality follows from the fact that $\varepsilon_{t,k}^{(i)}$ is sub-Gaussian with variance proxy in (E.3).
@@
(E.12) ${\mathbb{E}}_{{\mathcal{S}}}\Big[\sup_{b\in{\mathscr{B}}}{\mathcal{E}}_{T,R}(b)-2\widehat{{\mathcal{E}}}_{T,R,n}(b)\Big]\leq CR^{2}\sum_{k=1}^{d}\frac{\operatorname{VCdim}(\Pi_{k}{\mathscr{B}})}{n\log^{-1}(n)},$
@@
$\widehat{{\mathcal{E}}}_{T,R,n}^{k}(b)$ $=\frac{1}{n}\sum_{i=1}^{n}(b_{k}^{*}(t^{(i)},X_{t}^{(i)})-b_{k}(t^{(i)},X_{t}^{(i)}))^{2}\mathbbm{1}\{\|X_{t}^{(i)}\|_{\infty}\leq R\}.$
Applying Proposition 3.2, (E.1), and Lemma I.1 yields the following inequality
@@
which implies (E.12) by summing with respect to $1\leq k\leq d$.
@@
(E.13) ${\mathbb{E}}_{{\mathcal{S}}}\Big[\widehat{{\mathcal{E}}}_{T,R,n}(\widehat{b})\Big]\leq 2\inf_{b\in{\mathscr{B}}}{\mathcal{E}}_{T,R}(b)+C\lambda(T)R^{2}\Big(\sum_{k=1}^{d}\frac{\operatorname{VCdim}(\Pi_{k}{\mathscr{B}})}{n\log^{-1}n}+\frac{1}{\exp(\theta R^{2})}\Big),$
@@
[…]
$=\widehat{{\mathcal{E}}}_{T,R,n}(\widehat{b})+\frac{1}{n}\sum_{i=1}^{n}\sum_{k=1}^{d}(\varepsilon_{t,k}^{(i)})^{2}\mathbbm{1}\{\|X_{t}^{(i)}\|_{\infty}\leq R\}$
@@
Since $\widehat{b}$ is a minimizer of $\widehat{{\mathcal{L}}}_{T,n}(\cdot)$ over the hypothesis class $\mathscr{B}$}, it holds that $\widehat{{\mathcal{L}}}_{T,R,n}(\widehat{b})\leq\widehat{{\mathcal{L}}}_{T,n}(\widehat{b})\leq\widehat{{\mathcal{L}}}_{T,n}(b)$ for each $b\in{\mathscr{B}}$. Consequently,
[…]
@@
where we used the fact that ${\mathbb{E}}[\widehat{{\mathcal{L}}}_{T,n}(b)]={\mathcal{L}}_{T}(b)$ and ${\mathbb{E}}[\varepsilon_{t,k}^{(i)}b^{*}(t^{(i)},X_{t}^{(i)})]=0$. For the first term in the right-hand side of (E.14), we have
@@
(E.15) $={\mathcal{E}}_{T}(b)+{\mathbb{E}}_{{\mathcal{S}}}\Big[\frac{1}{n}\sum_{i=1}^{n}\sum_{k=1}^{d}(\varepsilon_{t,k}^{(i)})^{2}\mathbbm{1}\{\|X_{t}^{(i)}\|_{\infty}>R\}\Big]\leq{\mathcal{E}}_{T,R}(b)+2\frac{C\lambda(T)R^{2}}{\exp(\theta R^{2})},$
where the inequality follows from (E.9) and (E.11). It remains to bound the second term in the right-hand side of (E.14). Plugging Proposition 3.2 and (E.3) into Lemma I.2, we have
@@
(E.16) $\leq\frac{1}{4}{\mathbb{E}}_{{\mathcal{S}}}\Big[\widehat{{\mathcal{E}}}_{T,R,n}(\widehat{b})\Big]+C\lambda(T)R^{2}\sum_{k=1}^{d}\frac{\operatorname{VCdim}(\Pi_{k}{\mathscr{B}})}{n\log^{-1}n},$
where $C$ is a constant only depending $d$ and $\sigma$. Substituting (E.15) and (E.16) into (E.14) yields (E.13).
Finally, plugging (E.10), (E.12), and (E.13) into (E.4) completes the proof. ∎
Lemma E.2 (Approximation error).
@@
${\mathscr{B}}=\left\{b\in N(L,S):\begin{aligned} &\|b(t,x)\|_{\infty}\leq B_{\operatorname{vel}}R,\penalty 10000\ \|\partial_{t}b(t,x)\|_{2}\leq 3D\kappa(T)R,\\ &\|\nabla b(t,x)\|_{\operatorname{op}}\leq 3G,\penalty 10000\ (t,x)\in[0,T]\times{\mathbb{R}}^{d}\end{aligned}\right\},$
@@
Proof of Lemma E.2.
Denote by ${\mathbb{B}}_{R,T}^{\infty}=[0,T]\times{\mathbb{B}}_{R}^{\infty}$. According to Corollary J.6, for each element $1\leq k\leq d$, there exists a real-valued deep neural network $b_{k}$ with depth $L_{k}=\lceil\log_{2}(d+1)\rceil+3$ and number of parameters $S_{k}=(22(d+1)+6)(N+1)^{d+1}$, such that
@@
Set the hypothesis class as that in Lemma E.2. According to Lemma E.2, there exists a vector-valued deep neural network $b\in{\mathscr{B}}$ such that
(E.17) $\inf_{b\in{\mathscr{B}}}{\mathcal{E}}_{T,R}(b)\leq C\kappa^{2}(T)R^{2}N^{-2},$
where $C$ is a constant only depending on $d$ and $\sigma$. On the other hand, by applying Lemma B.12, the VC-dimension of this deep neural network class $\mathscr{B}$} is given as
(E.18) $\operatorname{VCdim}(\Pi_{k}{\mathscr{B}})\leq CN^{d+1}\log N,$
where $C$ is an absolute constant. Plugging (E.17) and (E.18) into Lemma E.1 yields
[…]
@@
(E.19) ${\mathbb{E}}_{{\mathcal{S}}}\big[{\mathcal{E}}_{T}(\widehat{b})\big]\leq C\kappa^{2}(T)R^{2}\Big(n^{-\frac{2}{d+3}}\log(n)+\exp(-\theta R^{2})\Big)$
@@
(E.20) $\,\mathrm{d}Z(t)$ $=b^{*}(t,Z(t))\,\mathrm{d}t,\quad t\in(0,T),$
@@
(E.21) $\,\mathrm{d}\widehat{Z}(t)$ $=\widehat{b}(t,\widehat{Z}(t))\,\mathrm{d}t,\quad t\in(0,T),$
@@
(E.22) $\leq 2\|Z(t)-\widehat{Z}(t)\|_{2}\|b^{*}(t,Z(t))-\widehat{b}(t,\widehat{Z}(t))\|_{2},$
where the second inequality follows from (E.20) and (E.21), and the inequality holds from Cauchy-Schwarz inequality. On the other hand, we find
(E.23) $\frac{\,\mathrm{d}}{\,\mathrm{d}t}\|Z(t)-\widehat{Z}(t)\|_{2}^{2}=2\|Z(t)-\widehat{Z}(t)\|_{2}\frac{\,\mathrm{d}}{\,\mathrm{d}t}\|Z(t)-\widehat{Z}(t)\|_{2}.$
Combining (E.22) and (E.23) implies
@@
Then we turn to estimate the distribution error using the particle error bound derived in Lemma E.3.
@@
According to Lemma E.3, we have
@@
where the first inequality follows from the definition of 22-Wasserstein distance and Jensen’s inequality. Substituting Theorem 3.10 into the above inequality completes the proof. ∎
Appendix F Proofs of Results in Section 3.4
@@
(F.1) $\widehat{Z}_{k}$ $=\widehat{Z}_{k-1}+\widehat{b}(t_{k-1},\widehat{Z}_{k-1})\tau,\quad 1\leq k\leq K,$
@@
Lemma F.1 (Discretization error of Euler method).
@@
(F.2) $=\widehat{Z}(t_{k})+\widehat{b}(t_{k},\widehat{Z}(t_{k}))\tau+\partial_{t}\widehat{b}(\theta,\widehat{Z}(\theta))\tau^{2},$
@@
(F.3) $\widehat{Z}_{k+1}=\widehat{Z}_{k}+\widehat{b}(t_{k},\widehat{Z}_{k})\tau.$
Subtracting (F.3) from (F.2) yields
@@
(F.4) $\leq(1+\|{\mathscr{B}}\|_{\operatorname{Lip}}\tau)\|\widehat{Z}(t_{k})-\widehat{Z}_{k}\|_{2}+D\kappa(T)R\tau^{2},$